OpenStudy (anonymous):

Simplify. Simplify: x^3y^5/x^2−5x−6*2x^2−15x+18/x^6y^5 I put 2x-3/x^3 but it's wrong can you please tell me the step by step process? Thank you

OpenStudy (anonymous):

\[\frac{ x ^{3}y ^{5} }{ x ^{2} }-5x-6*2x ^{2}-15x+\frac{ 18 }{ x ^{6} y ^{5}}\]

OpenStudy (anonymous):

@gg0824 right?

OpenStudy (anonymous):

\[\frac{ x^3y^5 }{ x^2-5x-6 }\times \frac{ 2x^2-15x+18 }{ x^6y^5 }\]

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

first can you factorize x^2-5x-6 ?

OpenStudy (anonymous):

\[\frac{ x^3y^5 }{ (x-6)(x+1) }\times \frac{ (2x-3)(x-6) }{ x^6y^5 } \]

OpenStudy (anonymous):

right then are u know if you have \[\frac{ A ^{b} }{ A ^{c} }=A ^{b-c}\]

OpenStudy (anonymous):

try to apply this